Equipment with a cost of $89,752, an estimated residual value of $6,499, and an estimated life of 13 years was depreciated by the straight-line method for 4 years. Due to obsolescence, it was determined that the remaining useful life should be shortened by 4 years and the residual value changed to zero. The depreciation expense for the current and future years is a. $25,616.31 b. $12,827.14 c. $4,933.51 d. $8,016.96
Answer #1

Annual Depreciation = (Cost - Salvage Value ) / Life in years

Annual Depreciation = (89,752-6499)/ 13 = $6,404

Book Value for the end of 4th year = 89,752 -(6404*4) = $64,136

Annual Depreication after REdcue the Life of the MAchine

Remaining life = 13-4-4 = 5 Years

= 64,136.69 / 5 Years = $12,827.14 (Answer)